# Challenges and strategies in verification of FastRICH ASIC for the LHCb RICH detector

### v1

@@ 1 Introduction
The LHCb RICH detector readout electronics and sensors will be replaced in LS3 Enhancements and II Upgrades II, respectively [1]. The FastRICH ASIC was designed to provide $25ps$ timing resolution, combined with continuous readout. The fine time resolution poses unique challenges to the design and verification. In particular, clock distribution, jitter, quantization error, and flip-flop metastability resolutions have major impact on the capability of the verification environment to assess the correctness of the ASIC output.
The common approaches used in verification, based on comparing the predicted with the effective output do not work. In this contribution, a novel approach based on approximate scoreboarding, is described. The verification framework was also re-used to run semi-directed tests, in form of scans, for verifying that the digital processing did not introduce artifacts or performance bottlenecks.
@@ 2 FastRICH ASIC
Figure 1: The FastRICH ASIC is composed of 16 channels, each with a dedicated digital processing. The variable-length data packets are processed, strictly in order, in the packet processor and frame builder. Before being streamed on the serial links the frames are encapsulated in the Aurora protocol.
The FastRICH ASIC, in figure 1, is a 16-channel readout ASIC with multiple configurable analogue front-ends in each channel, capable of resolving multiple hits per BX. A digital filtering (time gate) functionality allows configuring in which part of the BX the hits are recorded. In order to cope with the data rate variability, the ASIC implements a data-driven architecture with zero suppression. Moreover, the variable-length data format is optimized to minimize the output bandwidth. The data packets are encapsulated into the Aurora 64b-66b protocol [2], to ease data recovery in case of Single-Event Effects (SEEs). The data are transmitted over four configurable serial links ($320Mbps$ to $1.28Gbps$). The ASIC implements selective Triple-Modular Redundancy (TMR) to increase the SEE tolerance: the control logic is completely triplicated (combinatorial and sequential logic, resets, clocks, voters) whereas the data path is not triplicated.
@@ 3 Verification Strategy
In order to verify and validate the required timing resolution, the verification environment was carefully engineered to account, early-on, for the sources of uncertainty. The clock is provided to the ASIC with $5ps$ peak-peak jitter, which represents an upper bound of the jitter produced by an lpGBT elink [5, 6]; the PLL model is also accounting for a realistic amount of jitter, $8ps$ peak-peak, in line with worst case from analog simulations. We modeled metastability and metastability random resolution in selected flip-flops in the TDC already in RTL. This was a key step in verifying the functionality of the fully-digital TDC and validating its performance without the need of time-consuming implementation re-spins.
The protocol correctness is checked in all tests, under different conditions. In particular, it is to be noted that, depending on the channel occupancy, measured in hits per channel per BX, the ASIC can operate between two regimes:
  • Low occupancy. The ASIC mostly transmits empty packets. In this regime, the BX and Orbit tracking are the only piece of information which is transmitted. The verification framework should ensure that the timing information can be correctly tracked and that the latency introduced by the Aurora framing is within acceptable limits. No spurious packets should be present in this operating mode.
@@
  • Preparatory phase. In this phase the DUT is reset, randomly configured, then the PLL and fast control are locked. Finally, one hit is injected in each channel and it is readout. These hits act as a reference hit, used, later on, to calculate the expected time of each hit. During SEE verification, faults are injected in the DUT in this phase.
  • Repeated data taking phases. During these phases, the occupancy is randomized once for each repetition. Random hits are injected with the selected occupancy and they are readout. During the SEE verification, faults are injected in these phases. It is to be noted that, after each phase leading to data losses (for very high occupancy or SEEs), a subsequent phase with standard occupancy and no SEEs follows. The latter allows verifying that the data losses are only transitory and they do not require a reset of the DUT.
Additionally, we also employed formal checks to complement the functional verification framework: static and connectivity checks target at verifying that analog macros are correctly connected to the rest of the design. For the sake of brevity these will not be further treated in this contribution.
@@ 4 Verification of TDC timing performance
The random test described in section 3 was re-used to run semi-directed tests in forms of scans. These scans allowed looking at the ASIC behavior at a higher level of abstraction. Two examples developed for the FastRICH verification are the Time of Arrival (ToA) scan and Test Pulse (TP) scan. The first studies the TDC linearity, the latter the TP delivery network. In both cases one hit (or TP) injected in each BX (and each channel) each with an increasing ToA. The whole $25ns$ of a BX are swept with steps of $1ps$. These hits (or TPs) are readout and different tests are repeated with different time gate configurations. These scans allowed verifying that the TDC and clock distribution do not introduce artifacts in the data.
@@ 5 Verification of readout performance
Two additional type of scans were executed to study readout efficiency and readout latency vs data occupancy. These two scans allow verifying that the readout behaves consistently and predictably at different data occupancies. For both scans the test was repeated in different scenarios (number of output lanes enabled, data rate, etc.). The data from these scans were used to identify scenarios were the ASIC behavior was well within the specifications but it could be improved, providing important feedback to the designers.
Another important study executed was the validation of the TDC dead time. The digital TDC, should not limit the performance of the sensor or the analog front-end. In this scan two hits are injected in each BX at increasing distance in time $\delta_{t}$. The percentage of secondary hits is recorded vs $\delta_{t}$. An example, extracted from the gate-level simulation at typical corner is shown in figure 2.
Figure 2: TDC dead time scan. The requirement of being able to readout two hits in two consecutive BX translates in a TDC deadtime smaller than $18.87ns$, i.e. a $25ns$, a full BX, minus the maximum time gate duration, $6.25ns$. This corresponds to two hits in subsequent BXs, where the first hit is at the end of the time gate and the second hit is at the beginning of the time gate, in the subsequent BX. Here it can be observed that all the channels, in different colors, are well within the required limit and that they all behave consistently.
@@ 6 Conclusions
The verification strategy of the FastRICH ASIC was presented. The approach based on modeling metastability in key sequential elements in the digital TDC, together with an innovative scoreboarding technique, allowed verifying and validating the ASIC performance already at the RTL stage. This saved considerable amount of time in the iterations at the GLN verification. The presented approach re-used the tests, written for the constrained-random verification in order to study the behavior of the TDC at a higher level of abstraction, ensuring that no artifacts of the digital processing chain are present in the output data.

### v2

@@ 1 Introduction
The LHCb RICH detector readout electronics and sensors will be replaced in LS3 Enhancements and Upgrades II, respectively [1]. The FastRICH ASIC was designed to provide $25ps$ timing resolution, combined with continuous readout. The fine time resolution poses unique challenges to the design and verification. In particular, clock distribution, jitter, quantization error, and flip-flop metastability resolutions have a major impact on the capability of the verification environment to assess the correctness of the ASIC output.
The common approaches used in verification, based on comparing the predicted with the effective output, do not work. In this contribution, a novel approach based on approximate scoreboarding is described. The verification framework was also re-used to run semi-directed tests, in the form of scans, for verifying that the digital processing did not introduce artifacts or performance bottlenecks.
@@ 2 FastRICH ASIC
Figure 1: The FastRICH ASIC is composed of 16 channels, each with a dedicated digital processing. The variable-length data packets are processed, strictly in order, in the packet processor and frame builder. Before being streamed on the serial links, the frames are encapsulated in the Aurora protocol.
The FastRICH ASIC, in Figure 1, is a 16-channel readout ASIC with multiple configurable analogue front-ends in each channel, capable of resolving multiple hits per BX. A digital filtering (time gate) functionality allows rejecting hits that occur outside a configurable window inside each BX. In order to cope with the data rate variability, the ASIC implements a data-driven architecture with zero suppression. Moreover, the variable-length data format is optimized to minimize the output bandwidth. The data packets are encapsulated into the Aurora 64b-66b protocol [2], to ease data recovery in case of Single-Event Effects (SEEs). The data are transmitted over four configurable serial links ($320Mbps$ to $1.28Gbps$). The ASIC implements selective Triple-Modular Redundancy (TMR) to increase the SEE tolerance: the control logic is completely triplicated (combinatorial and sequential logic, resets, clocks, voters), whereas the data path is not triplicated.
@@ 3 Verification Strategy
In order to verify and validate the required timing resolution, the verification environment was carefully engineered to account, early-on, for the sources of uncertainty. The clock is provided to the ASIC with $5ps$ peak-peak jitter, which represents an upper bound of the jitter produced by an lpGBT elink [5, 6]; the PLL model also accounts for a realistic amount of jitter, $8ps$ peak-peak, in line with worst case from analog simulations. We modeled metastability and metastability random resolution in selected flip-flops in the TDC already in RTL. This was a key step in verifying the functionality of the fully-digital TDC and validating its performance without the need for time-consuming implementation re-spins.
The protocol correctness is checked in all tests under different conditions. In particular, it is to be noted that, depending on the channel occupancy, measured in hits per channel per BX, the ASIC can operate between two regimes:
  • Low occupancy. The ASIC mostly transmits empty packets. In this regime, the LHC BX and Orbit [7] tracking [^2] BXID wraps around at every LHC Orbit. Orbit tracking refers to keeping track of which LHC Orbit a specific BXID corresponds to. are the only pieces of information that are transmitted. The verification framework should ensure that the timing information can be correctly tracked and that the latency introduced by the Aurora framing is within acceptable limits. No spurious packets should be present in this operating mode.
@@
  • Preparatory phase. In this phase, the DUT is reset, randomly configured, and then the PLL and fast control are locked. Finally, one hit is injected in each channel and it is readout. These hits act as a reference hit, used later on, to calculate the expected time of each hit. During SEE verification, faults are injected in the DUT in this phase.
  • Repeated data taking phases. During these phases, the occupancy is randomized once for each repetition. Random hits are injected with the selected occupancy and they are readout. During the SEE verification, faults are injected in these phases. It is to be noted that, after each phase leading to data losses (for very high occupancy or SEEs), a subsequent phase with standard occupancy and no SEEs follows. The latter allows verifying that the data losses are only transitory and that they do not require a reset of the DUT.
Additionally, we also employed formal checks to complement the functional verification framework: static and connectivity checks target at verifying that analog macros are correctly connected to the rest of the design. For the sake of brevity, these will not be further treated in this contribution.
@@ 4 Verification of TDC timing performance
The random test described in section 3 was re-used to run semi-directed tests in the form of scans. These scans allowed looking at the ASIC behavior at a higher level of abstraction. Two examples developed for the FastRICH verification are the Time of Arrival (ToA) scan and Test Pulse (TP) scan. The first studies the TDC linearity, the latter the TP delivery network. In both cases, one hit (or TP) was injected in each BX (and each channel) each with an increasing ToA. The whole $25ns$ of a BX are swept with steps of $1ps$. These hits (or TPs) are readout and different tests are repeated with different time gate configurations. These scans allowed verifying that the TDC and clock distribution do not introduce artifacts in the data.
@@ 5 Verification of readout performance
Two additional types of scans were executed to study readout efficiency and readout latency vs data occupancy. These two scans allow verifying that the readout behaves consistently and predictably at different data occupancies. For both scans, the test was repeated in different scenarios (number of output lanes enabled, data rate, etc.). The data from these scans were used to identify scenarios where the ASIC behavior was well within the specifications but could be improved, providing important feedback to the designers.
Another important study executed was the validation of the TDC dead time. The digital TDC, should not limit the performance of the sensor or the analog front-end. In this scan, two hits are injected in each BX at increasing distance in time $\delta_{t}$. The percentage of secondary hits is recorded vs $\delta_{t}$. An example, extracted from the gate-level simulation at typical corner is shown in Figure 2.
Figure 2: TDC dead time scan. The requirement of being able to readout two hits in two consecutive BX translates to a TDC deadtime smaller than $18.87ns$, i.e. a $25ns$, a full BX, minus the maximum time gate duration, $6.25ns$. This corresponds to two hits in subsequent BXs, where the first hit is at the end of the time gate and the second hit is at the beginning of the time gate, in the subsequent BX. Here, it can be observed that all the channels, in different colors, are well within the required limit and that they all behave consistently.
@@ 6 Conclusions
The verification strategy of the FastRICH ASIC was presented. The approach based on modeling metastability in key sequential elements in the digital TDC, together with an innovative scoreboarding technique, allowed verifying and validating the ASIC performance already at the RTL stage. This saved a considerable amount of time in the iterations at the GLN verification. The presented approach re-used the tests written for the constrained-random verification in order to study the behavior of the TDC at a higher level of abstraction, ensuring that no artifacts of the digital processing chain are present in the output data.